OpenStudy (anonymous):

@sleepyjess @satellite73 @Leader @perl @pooja195

OpenStudy (anonymous):

Kylie sells 2 concert tickets every minute. She has already sold 8 tickets. If f(t) represents the total number of tickets that Kylie will sell in t minutes, which of the following functions represents the relationship between f(t) and t? f(t) = 8t - 2 f(t) = 2t - 8 f(t) = 2t + 8 f(t) = 8t + 2

OpenStudy (anonymous):

c

OpenStudy (apoorvk):

What makes you think C? :P

OpenStudy (apoorvk):

It is right btw, because 8 tickets have been sold already. So, '8' can be a constant here. And she sells '2' times 't' tickets, by the time 't' minutes have passed. So, she sells (2 x t) + 8 tickets, (since 8 had been sold already).

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1 fourthx + 2. y = 4x - 1 y = 4x - 3 y = 1 fourthx - 3 y = 1 fourthx - 1

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

lol i read it wrong try using numbers, not words that is why mathematicians invented things like \(\frac{1}{4}\)

OpenStudy (anonymous):

ok well am i correct ?

OpenStudy (anonymous):

the slope is \(\frac{1}{4}\) for sure how about the y intercept?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

not for the line with slope \(\frac{1}{4}\) through \((-4,-2)\)

OpenStudy (anonymous):

put \[y+2=\frac{1}{4}(x+4)\] and solve for \(y\)

OpenStudy (anonymous):

may you help me w, that

OpenStudy (apoorvk):

Do you know this: Two lines parallel will have the same slope. And. The value of the slope comes from \(m\), comparing with this standard equation: \[y = mx + c\] Rest you substitute the values of \(a,b\), that is coordinates of point through which the line is passing and find out the constant intercept, \(c\). After that, just replace the value of 'm' and 'c' and you have it. Did you follow this?

OpenStudy (anonymous):

y =1/4x -1 ? @satellite73

OpenStudy (anonymous):

\[y+2=\frac{1}{4}(x+4)\] solve for y 2 steps always the same 1) distribute \[y+2=\frac{1}{4}x+1\]

OpenStudy (anonymous):

2) solve for \(y\)\[y=\frac{1}{4}x-1\] what you said

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

please help me w. that xd

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

y=-4x

OpenStudy (anonymous):

@satellite73